OpenStudy (anonymous):

Can i get help with this problem dealing with finding the area between curves: x=y^2-y and x=2y^2-2y-6

OpenStudy (anonymous):

I hope youre taking calculus. Take the integrals of both, then subtract the line thats underneath from the one thats over it.

OpenStudy (anonymous):

i dont quite get you. What line?

OpenStudy (anonymous):

what line do i subtract?

OpenStudy (anonymous):

Youre given two equations which can be graphed. One line will be above the other on the graph. Once you take the integrals of both, you now will have the areas of both curves. to find the areas between them, subtract the smaller one from the larger one.

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/CalcI/AreaBetweenCurves.aspx Go to example three

OpenStudy (anonymous):

okay am i integrating with respects to y?

OpenStudy (anonymous):

it might help if you solve for y first too.

OpenStudy (anonymous):

oh crap. you might have to do implicit differentiation first. have you done that?

OpenStudy (anonymous):

oh so solve for y in both equations

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

solve for y, then integrate.

OpenStudy (anonymous):

i still dont understand how to get started

OpenStudy (anonymous):

yeah we can do this one quickly

OpenStudy (anonymous):

first of all i hate this x = stuff so lets make it \[y=x^2-x\] and \[y=2x^2-2x-6\]

OpenStudy (anonymous):

we are going to get a number out of this, so the variable is irrelevant

OpenStudy (anonymous):

we need only the limits of integration, which is where these two parabolas meet

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

set them equal and solve

OpenStudy (anonymous):

and thats it, nothing special?

OpenStudy (anonymous):

like resolve for y or somehting like that?

OpenStudy (anonymous):

i get x = 0 and x = 2

OpenStudy (anonymous):

then you have to integrate the bigger one minus the smaller one over that interval

OpenStudy (anonymous):

now, which one would be the smaller one and bigger one?

OpenStudy (anonymous):

damn damn damn i messed up

OpenStudy (anonymous):

lets go back to \[x=y^2-y\] \[ x=2y^2-2y-6\]

OpenStudy (anonymous):

lets match these up

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

set them equal and solve

OpenStudy (anonymous):

\[2y^2-2y-6=y^2-y\] \[y^2-y-6=0\] \[(y-3)(y+2)=0\] \[y=-2,y=3\]

OpenStudy (anonymous):

so those are our limits of integration from -2 to 3

OpenStudy (anonymous):

we are going to integrate along the y axis, that is wrt y. no different than wrt x, just dy instead of dx

OpenStudy (anonymous):

but how come, the curves are top and bottome, i thought that when integrating with respects to y there should be a right and left curve

OpenStudy (anonymous):

when i graphed them i saw that they were like two parabolas one on top and one on bottome

OpenStudy (anonymous):

there is

OpenStudy (anonymous):

we go from y = -2 to y = 3

OpenStudy (anonymous):

lets graph

OpenStudy (anonymous):

which is left curve will be f and which will be g

OpenStudy (anonymous):

it will be \[y^2-y-(2y^2-2y-6)\] \[-y^2+y+6\] from -2 to 3

OpenStudy (anonymous):

\[\int_{-2}^3(-y^2+y+6)dy\]

OpenStudy (anonymous):

ooooooooooooh i have an idea. lets see if it works!

OpenStudy (anonymous):

lets cheat and see if we can get it instantly. i know it will not explain, but we have it now in any case

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

all you have to do is ask! http://www.wolframalpha.com/input/?i=area+between+the+curves+x%3Dy^2-y+and+x%3D2y^2-2y-6

OpenStudy (anonymous):

with the method you have shown me, do i have to compensate in some way for the portion below the x-axis

OpenStudy (anonymous):

yeah i screwed that up so forget about it. it was just plain wrong

OpenStudy (anonymous):

oh i mean the switching x and y was wrong. the method above is correct

OpenStudy (anonymous):

oh man , i need to do this for HW

OpenStudy (anonymous):

oh, so then the other way is still good

OpenStudy (anonymous):

you don't have to worry about "above" and "below" because you are taking one function minus another

OpenStudy (anonymous):

second way is good. switching was stupid

OpenStudy (anonymous):

so then if i evluate the integral of -y^2+y+6 at -2 and 3 i should get my answer right

OpenStudy (anonymous):

here look at this here is your integral not as an area between two curves. you see that everything is positive http://www.wolframalpha.com/input/?i=integrate+-2+to+3++%28-y^2%2By%2B6%29+dy

OpenStudy (anonymous):

okay so then just integrat that and then evaluate right

OpenStudy (anonymous):

yes, find the anti-derivative, plug in 3, plug in -2, subtract, get \[\frac{125}{3}\] and be done

OpenStudy (anonymous):

yup you got it!

OpenStudy (anonymous):

thanks man, i appreaciate it

OpenStudy (anonymous):

yw